OpenStudy (anonymous):

36 candy are in a bag 4/9 have been eaten how many have been eaten?

OpenStudy (anonymous):

mmm candy \[36\times \frac{4}{9}=\frac{36\times 4}{9}=4\times 4=16\]

OpenStudy (anonymous):

could somby help w mine ?